I have a word problem with ages and I need to know what euaation I would use to solve it. First you should know the ages in question are 27 and 8.here is the question: At some point during the lives of you(27) and the youngest person(8), your age will be three times his/her age at that moment. Write an equation which models how old in years each of you will be when you are three times as old as the younger person.
Answer by Theo(13342)

let x be the number of years from now.
27 + x = 3 * (8 + x)
multiply out to get:
27 + x = 24 + 3x
subtract x from both sides of the equation to get:
27 = 24 + 2x
subtract 24 from both sides of the equation to get:
2x = 3
divide both sides of the equation by 2 to get:
x = 1.5
in 1.5 years, your age will be 28.5 years old and the younger person's age will be 9.5
3 * 9.5 = 28.5